### v1

# Sign-Problem-Free Nuclear Quantum Monte Carlo

Abstract
Quantum Monte Carlo (QMC) methods offer exact solutions for quantum many-body systems but face severe limitations in fermionic systems like atomic nuclei due to the sign problem. While sign-problem-free QMC algorithms exist, they have been confined to simple models with limited predictive power for real nuclei. Here we overcome this barrier by developing a novel lattice nuclear force that is rigorously sign-problem-free for even-even nuclei. This interaction achieves a standard deviation of $\sigma=2.932$ MeV from experimental binding energies for 76 even-even nuclei ($N,Z\leq 28$), matching state-of-the-art phenomenological mean-field models. Key innovations include the first sign-problem-free implementation of spin-orbit coupling for shell evolutions and an efficient QMC-optimized framework for global parameter fitting. Using this approach, we compute binding energies from 4He to 132Sn, symmetric nuclear matter saturation, and reveal novel spin-orbit-driven clustering in light nuclei. This work transforms sign-problem-free QMC into a scalable and predictive nuclear structure tool, establishing a non-perturbative foundation for ab initio calculations extending to heavy nuclei.
Introduction
Understanding the emergence of the nuclear structure from fundamental nucleon-nucleon interactions is a central challenge in modern nuclear physics. While exact diagonalization approaches are limited by the exponential growth of Hilbert space dimensionality, quantum Monte Carlo (QMC) methods offer a promising solution by stochastically sampling the vast Hilbert space, reducing computational complexity to polynomial scaling with nucleon number. Recent years have seen numerous QMC applications to nuclear many-body problems (see Refs. PhysRept278-1 ; PPNP47-319 ; PPNP63-117 ; Lahde2019nuclear ; RMP87-1067 ; AnnRevNuclPartSci69-279 ; RMP94-031003 for reviews).
Despite significant successes, most QMC applications are fundamentally constrained by the fermionic sign problem PRL94-170201 . This arises when integrating highly oscillatory functions, where cancellations between positive and negative amplitudes lead to exponentially large statistical noise scaling with particle number. Although mitigation techniques exist, e.g., fixed-node JChemPhys77-349 ; JChemPhys77-5593 ; IntlRevPhysChem14-85 /constrained path approximations PRL74-3652 ; PRB55-7464 ; PRC62-014001 , the Lefschetz thimble method PRD86-074506 ; PRD88-051501R , and perturbative expansions PRL128-242501 ; PRR5-L042021 ; EPJA61-85 ; PRC111-015801 , their implementations present significant numerical challenges and often introduce systematic biases.
Notably, for certain symmetry-protected fermionic systems, QMC calculations are immune to the sign problem. In these rare cases, QMC delivers unbiased results with high precision and exceptional scalability. For example, in lattice quantum chromodynamics (QCD), the isospin symmetry between the up and down quarks implies that the fermionic determinant factorizes into a perfect square, guaranteeing positive amplitudes for any gauge-field configuration ProgTheorPhys110-615 . It was further shown that a broader class of interactions without a factorizable determinant can also be sign-problem-free, where positivity is protected by the time-reversal symmetry of the fermion matrix PRB71-155115 ; PRB91-241117 . See Ref. AnnuRevCondMattPhys10-337 and references therein for sign-problem-free QMC applied to phenomenological condensed matter models.
In nuclear physics, however, sign-problem-free QMC algorithms remain underexploited, particularly for ab initio approaches. Such calculations employ complicated high-fidelity nuclear forces derived from fundamental symmetries, unavoidably inducing sign problems PRL111-032501 ; PRL113-192501 ; PRC98-044002 ; Nature630-59 ; PRL132-232502 . Critically, any sign problem, even negligible for light nuclei, grows exponentially with particle number, preventing direct QMC simulations of heavy nuclei. Remedies like perturbation theory PRL128-242501 ; PRR5-L042021 require a nearly sign-problem-free interaction as the non-perturbative zeroth order. Meanwhile, recent ab initio calculations found inconsistencies among few-body, many-body, and nuclear matter predictions PLB736-119 ; PRC104-064312 ; PRC104-054001 ; FewBodySyst64-77 , which stimulates the development of minimal interactions attempting to reconcile these contradictory results PLB797-134863 ; PRL121-072701 ; PRL133-142501 . These minimal interactions typically predict nuclear observables with percent-level accuracy and are potentially solvable using sign-problem-free QMC because of their simplicity. An example is the sign-problem-free Wigner-SU(4) force PLB797-134863 ; PhysRev51-106 ; PRL117-132501 ; PRL127-062501 that successfully reproduces light-nuclei binding energies and charge radii, and describes emergent geometry in 12NatComm14-2777 , clustering in hot nuclear matter PRL125-192502 ; PLB850-138463 , Beryllium isotope spectra PRL134-162503 , and $\alpha$-particle monopole transition PRL132-062501 . These achievements motivate subsequent searches for more accurate sign-problem-free interactions. Here, the core challenge is that time-reversal symmetry enabling sign-problem-free simulations forbids many essential interactions like the tensor force and most spin/isospin-dependent terms PRL95-232502 ; PPNP76-76 ; PRL89-182501 ; PRL98-132501 , severely constraining the available model space.
In this Letter, we tackle this challenge by constructing a novel minimal nuclear force that is simple enough to avoid the Monte Carlo sign problem, yet sufficiently sophisticated to reproduce key experimental data. To satisfy these competing requirements, we introduce a sign-problem-free spin-orbit interaction in addition to the Wigner-SU(4) force to capture shell evolution, and optimize its parameters against selected nuclear binding energies using an efficient derivative-based method. While conceptually inspired by phenomenological mean-field force construction RMP75-121 ; PPNP57-470 ; JPG46-013001 , crucially, our approach yields exact solutions to the Hamiltonian.
Method
We formulate the QMC algorithm within the framework of the ab initio approach of nuclear lattice effective field theory (NLEFT), where the nuclear force models are discretized and numerically solved on a three-dimensional cubic lattice using auxiliary field Monte Carlo methods. Recent advances in NLEFT have yielded significant achievements across multiple domains, e.g., nuclear ground states EPJA31-105 ; PRL104-142501 ; EPJA45-335 ; PLB732-110 and excited states PRL112-102501 , intrinsic density distributions and clustering phenomena PRL106-192501 ; PRL109-252501 ; PRL110-112502 ; PRL119-222505 ; arXiv2411.17462 , nucleus-nucleus scattering dynamics PRC86-034003 ; Nature528-111 , and hypernucleus PRL115-185301 ; EPJA56-24 ; EPJA60-215 ; PRD111-036002 .
The minimal nuclear force employed here comprises both two- and three-body interactions supplemented by a spin-orbit coupling term. The Hamiltonian is defined on a $L^{3}$ cubic lattice with integer coordinates $\bm{n}=(n_{x},n_{y},n_{z})$,
$H=\sum_{\bm{n}}\left[-\frac{\Psi^{\dagger}\nabla^{2}\Psi}{2M}+:\frac{C_{2}}{2}\overline{\rho}^{2}+\frac{C_{3}}{6}\overline{\rho}^{3}+C_{s}\overline{\rho}\ \overline{\rho}_{s}:\right],$ (1)
where the summation spans all lattice sites and the colons mean the normal ordering. Here $\Psi(\bm{n})$ and $\Psi^{\dagger}(\bm{n})$ denote annihilation and creation operators, $M=938.92$ MeV is the nucleon mass, $\nabla^{2}$ is the Laplace operator implemented via fast Fourier transform (FFT), and $C_{2}$, $C_{3}$, $C_{s}$ are coupling constants. We fix the lattice spacing to $a=1.32$ fm throughout and use symbols with overline to denote smeared quantities,
$\overline{\sigma}(\bm{n})=\sigma(\bm{n})+s_{\rm L}\sum_{|\bm{n}^{\prime}-\bm{n}|=1}\sigma(\bm{n}^{\prime}),\qquad\sigma=\rho,\rho_{s},$ (2)
with $s_{\rm L}$ a real number controlling local smearing. The densities $\rho$ and $\rho_{s}$ are defined as
$\rho$ $=$ $\overline{\Psi}^{\dagger}\overline{\Psi},$
$\rho_{s}$ $=$ $\frac{i}{4}\sum_{ijk}\epsilon_{ijk}\nabla_{i}\left[\overline{\Psi}^{\dagger}(\overrightarrow{\nabla_{j}}-\overleftarrow{\nabla_{j}})\sigma_{k}\overline{\Psi}\right],$ (3)
where $\epsilon_{ijk}$ is the Levi-Civita symbol, $\nabla$ represents the spatial gradients implemented with the FFT, and $\sigma_{k}$ is the spin Pauli matrices. The smeared operators $\overline{\Psi}$ and $\overline{\Psi}^{\dagger}$ are defined as,
$\overline{\sigma}(\bm{n})=\sigma(\bm{n})+s_{\rm NL}\sum_{|\bm{n}^{\prime}-\bm{n}|=1}\sigma(\bm{n}^{\prime}),\qquad\sigma=\Psi,\Psi^{\dagger}$ (4)
with $s_{\rm NL}$ a real number controlling non-local smearing. Implicit spin/isospin indices in $\Psi(\bm{n})$, $\Psi(\bm{n})^{\dagger}$, and $\sigma_{k}$ are summed. Both local (Eq. (2)) and nonlocal (Eq. (4)) smearing govern the effective interaction range. The relative strength between $s_{\rm L}$ and $s_{\rm NL}$ determines the locality, an essential element for nuclear binding. In particular, the heavy nuclei dissociate into small clusters ($A\leq 4$) at the non-local limit ($s_{\rm L}=0$) and exhibit severe overbinding at the local limit ($s_{\rm NL}=0$PRL117-132501 . The three-body force ($C_{3}$) significantly improves the light nuclei structure PLB797-134863 . The newly introduced spin-orbit term ($C_{s}$) generates requisite shell structure evolution. For $C_{s}=0$ we return to the Wigner-SU(4) interaction thoroughly studied in Ref. PLB797-134863 . We also incorporate a perturbative Coulomb force for protons SM .
We solve the Hamiltonian Eq. (1) using the imaginary time projection method,
$|\Phi\rangle\propto\lim_{\tau\rightarrow\infty}\exp(-\tau H)|\Phi_{T}\rangle,$ (5)
where $|\Phi\rangle$ is the ground state and $|\Phi\rangle_{T}$ is a trial wave function with non-zero overlap. For sufficiently large $\tau$, excited state contributions are exponentially suppressed. Expectation values are computed through numerical extrapolation
$\langle O\rangle=\lim_{\tau\rightarrow\infty}\frac{\langle\Phi_{T}|e^{-\tau H/2}Oe^{-\tau H/2}|\Phi_{T}\rangle}{\langle\Phi_{T}|e^{-\tau H}|\Phi_{T}\rangle}.$ (6)
In NLEFT we implement Eq. (6) via auxiliary field transformation. For the two-body forces,
$e^{-a_{t}H}$ $=$ $\int\mathcal{D}c\exp\left[-\sum_{\bm{n}}\frac{c(\bm{n})^{2}}{2}+\frac{a_{t}}{2M}\Psi^{\dagger}\nabla^{2}\Psi\right.$ (7)
$\left.+\sqrt{-a_{t}C_{2}}\sum_{\bm{n}}c(\bm{n})\left(\overline{\rho}(\bm{n})+\frac{C_{s}}{2C_{2}}\overline{\rho}_{s}(\bm{n})\right)\right],$
where $a_{t}$ is the temporal step and $c(\bm{n})$ is a real auxiliary field. The terms of order $O(C_{s}^{2})$ generated from this transformation are negligible. (Transformation involving 3NF follows SM ). The $c$ fields at each imaginary time step are collectively integrated using the importance sampling Monte Carlo. Without loss of generality, we consider even-even nuclei and take the trial wave function $|\Phi_{T}\rangle$ as a Slater determinant of nucleon wave functions paired by time-reversal symmetry $\mathcal{T}=i\sigma_{y}\mathcal{K}$, where $\mathcal{K}$ is the complex conjugate. These nucleons interact with the $c$-field and propagate independently under Eq. (7). As the single-particle Hamiltonian in Eq. (7) commute with $\mathcal{T}$, the nucleons remain paired during the imaginary-time evolution. The resulting determinant of the fermionic correlation matrix is positively definite, ensuring completely sign-problem-free simulations. (See SM for proof.)
The interaction in Eq. (1) contains five parameters $C_{2}$, $C_{3}$, $C_{s}$, $s_{\rm L}$, and $s_{\rm NL}$. We found that the mass prediction is not sensitive to $s_{\rm NL}$ and fix $s_{\rm NL}=0.45$ throughout. The remaining four parameters are optimized against experimental binding energies of 4He, 16O, 24Mg, 28Si, 32S and 40Ca. Here we select three doubly-magic nuclei and three open-shell nuclei to constrain both the bulk properties and the shell evolution. The optimization minimize the loss function
$\chi^{2}=\sum_{A}\left[\frac{E(A)-E_{{\rm exp}}(A)}{\Delta(A)}\right]^{2},$ (8)
with weights $\Delta(A)=1$ MeV for 4He and $3$ MeV for other nuclei. To apply the steepest descent method, we need to calculate the partial derivatives of $\chi^{2}$ against the fitting parameters, which in turn depend on the derivatives of the binding energies $E(A)$. The latter was computed using the Feynman-Hellmann theorem,
$\partial E(A)/\partial x=\langle\Phi|H(x+\delta)-H(x)|\Phi\rangle/\delta,$ (9)
where $x$ is one of the fitting parameters and $\delta$ is a small number chosen for convenience. The expectation values in Eq. (9) are measured concurrently with other observables in Monte Carlo simulations. Crucially, our sign-problem-free QMC enables precise, unbiased estimation of all derivatives simultaneously, facilitating rapid convergence with adaptive learning rate. The optimization with a reasonable initial guess typically converges within ten iterations. Among multiple local minima, we identify the globally optimal values $C_{2}=-4.410\times 10^{-7}$ MeV-2, $C_{3}=1.561\times 10^{-15}$ MeV-5, $C_{s}=8.590\times 10^{-12}$ MeV-4 and $s_{\rm L}=8.082\times 10^{-2}$. In what follows we denote this group of parameters as LAT-OPT1. Our strategy, distinct from derivative-free methods used in recent chiral force optimizations against finite nuclei PRL110-192502 ; PLB761-87 , exhibits an efficiency independent of parameter space dimensionality due to simultaneous derivative measurement.
Fig. 1: Ground state energies from LAT-OPT1 (Errors smaller than the symbols). Red circles enclose the target nuclei for calibrating the interaction. Crosses denote experimental values exp . Isotopic chains are connected to guide the eyes. See SM for numerical results.
Results
[…]
In Fig. 1 we show binding energies for $76$ even-even nuclei with $N,Z\leq 28$. The results show excellent agreement with experimental values, yielding a standard error of $\sigma=2.932$ MeV. For comparison, state-of-the-art density functional theories typically give $\sigma\approx 2$-$3$ MeV in this mass region PRC85-024304 ; mean_field_mass_tables ; PRC104-054312 . Note that our model achieves this level of accuracy with only four parameters trained on six nuclei. This remarkable generalization capability can be attributed to the incorporation of full quantum correlations, with shell effect, pairing, clustering, deformation, collective rotation, and shape fluctuation all treated on an equal footing. Systematic overbinding at large neutron numbers indicates missing components in our nuclear force, such as the pion-exchange potentials dictated by the chiral dynamics RMP81-1773 ; PR503-1 ; frontierinphysics8-98 ; arXiv:2402.14032 .
[…]
While existing nuclear QMC approaches using high-fidelity forces remain confined to light/medium-mass systems and require error mitigation, our sign-problem-free method can be directly used to investigate heavy nuclei. In Fig. 2 we plot the QMC results for 100Sn, extrapolated to infinite imaginary time. The statistical errors are sufficiently small to yield total uncertainties below $1$ MeV. We also probe dependencies on temporal step size and box volume, finding only negligible variations. Moreover, The calculated binding energy agrees with experiment within $2$ MeV. These remarkable one-thousandth level computational and theoretical accuracies demonstrate sign-problem-free QMC as a scalable, high-precision approach for heavy nuclei.
Tab. 1 lists calculated binding energies for representative nuclei. The relative errors against the experiments are below 3% for 4He and 132Sn, and below 1% for other nuclei, demonstrating high accuracy across the nuclear chart. Furthermore, to quantify the impact of the spin-orbit coupling term proportional to $C_{s}$, we also list its partial contributions $E_{\rm sl}$. This term accounts for 5-10% of the binding energies on average and systematically fluctuates with the number of nucleons. Notably, the fraction $E_{\rm sl}/E_{\rm bind}$ is always below 5% for nuclei with traditional harmonic oscillator magic numbers 2, 8, 20, 40 such as 4He, 16O, 40Ca and 80Zr, but significantly enhanced for nuclei with new magic numbers 28, 50, 82 such as 56Ni, 100Sn and 132Sn. This extra binding reflects the enhanced stability of these nuclei obtained from shell gaps generated by spin-orbit splitting, directly validating the Mayer-Jensen shell model paradigm within our framework without employing a single-particle mean field.
Table 1: Calculated binding energies $E_{\rm bind}$, expectation values of the spin-orbit coupling term $E_{\rm sl}$, their ratio and experimental binding energies exp (EXP) for selected nuclei.
Nucleus $E_{\mathrm{bind}}$ (MeV) $E_{\mathrm{sl}}$ (MeV) $E_{\mathrm{sl}}/E_{\mathrm{bind}}$ EXP (MeV)
4He $-29.0(2)$ $-0.3$ 0.010 $-28.3$
12C $-91.3(1)$ $-13.3$ 0.146 $-92.2$
14C $-104.6(1)$ $-12.7$ 0.121 $-105.3$
16O $-126.9(2)$ $-5.6$ 0.044 $-127.6$
40Ca $-343.0(2)$ $-13.6$ 0.040 $-342.1$
48Ca $-414.5(3)$ $-42.3$ 0.102 $-416.0$
56Ni $-479.3(6)$ $-74.6$ 0.156 $-484.0$
80Zr $-672.1(8)$ $-23.3$ 0.035 $-669.2$
90Zr $-782.1(5)$ $-64.8$ 0.083 $-783.9$
100Sn $-824.7(8)$ $-103.0$ 0.125 $-825.2$
132Sn $-1134.2(27)$ $-110.9$ 0.098 $-1102.8$
Our approach solve the quantum many-body problem non-perturbatively, enabling studies of strong-correlation phenomena like nuclear clustering that emerges from the intricate interplay among shell structure, quantum fluctuations and continuum effects PhysRept432-43 ; RMP90-035004 . Fig. 3 (a) shows the loss function $\chi^{2}$ in Eq. (8) as a function of the spin-orbit coupling constant $C_{s}$, with other parameters fixed to LAT-OPT1 values. A pronounced minimum occurs precisely at the LAT-OPT1 value $C_{s}^{\rm opt}$ (vertical line), confirming that our optimization robustly identifies the physical spin-orbit strength stringently constrained by target nuclei. Fig. 3 (b) displays binding energies for 4He and light $\alpha$-conjugate nuclei (8Be, 12C and 16O) relative to $n$-$\alpha$ breakup thresholds. The 4He binding is independent of $C_{s}$, while 8Be persists unbound against two-$\alpha$ decay. The 3-$\alpha$ (12C) and 4-$\alpha$ (16O) systems are more sensitive to spin-orbit coupling. In the SU(4)-symmetric limit ($C_{s}\to 0)$, 16O is stable, but 12C is unbound against $3\alpha$ dissociation. Conversely, at the other extreme of strong spin-orbit coupling ($C_{s}\gg C_{s}^{\rm opt}$), the 3-$\alpha$ and 4-$\alpha$ binding energies coincide, indicating that 12C becomes deeply bound, while 16O turns unstable against 12C$+\alpha$ breakup. Remarkably, only near the $\chi^{2}$-optimized value $C_{s}^{\rm opt}$ do we reproduce the experimental binding hierarchy. This reveals a direct connection between spin-orbit coupling and effective $\alpha$-cluster interactions, suggesting a unified mechanism governing shell evolution and clustering in light nuclei.
Figure 3: (a) Loss function (Eq. (8)) versus spin-orbit coupling $C_{s}$. (b) Binding energies relative to $n$-$\alpha$ thresholds: 8Be (squares), 12C (circles), 16O (diamonds). Full symbols: experimental values exp ; solid/open triangles: calculated/experimental 4He energies. Vertical line: optimal $C_{s}^{\rm opt}$ minimizing $\chi^{2}$.
Finally, Fig. 4 shows the predicted equation of state for symmetric nuclear matter. Comparison to the Wigner-SU(4) interaction PLB797-134863 reveals significant improvement. We identify a saturation point at $(\rho_{0},E/A)=(0.175$ fm-3, $-15.4$ MeV)))), consistent with empirical values.
Figure 4: Equation of state for symmetric nuclear matter calculated with LAT-OPT1 (circles) and Wigner-SU(4) PLB797-134863 (diamonds) interactions. Full symbols denote corresponding saturation points. Green rectangle mark the empirical saturation point (0.16±0.01(0.16$\pm$ 0.01( 0.16 ± 0.01 fm-3, $-16.0\pm 1.0$ MeV)))) Bethe1971 .
Summary and perspective
We have developed a sign-problem-free lattice nuclear force, enabling unbiased QMC simulations across the nuclear chart. This interaction reasonably reproduces experimental binding energies from 4He to 132Sn and nuclear matter saturation with statistical errors of around $1$ MeV at a feasible computational cost. The approach can be directly extended to investigate excited-state spectra PRL106-192501 , density correlations PRL119-222505 , and finite-temperature nuclear properties PRL125-192502 using advanced lattice QMC algorithms. Future refinements will incorporate more nuclei and additional nuclear observables to further constrain the interaction. Moreover, this interaction provides an robust non-perturbative foundation for ab initio calculations with high-fidelity nuclear chiral forces. The missing physics in our model can be incorporated order-by-order perturbatively PRL128-242501 ; Nature630-59 . Finally, this work provides a scalable and systematically improvable computational paradigm that bridges ab initio calculations and phenomenological models.
Acknowledgement
We thank members of the Nuclear Lattice Effective Field Theory Collaboration for insightful discussions, and Dean Lee for critical reading of the manuscript. This work has been supported by NSAF No. U2330401 and National Natural Science Foundation of China with Grant No. 12275259.
Supplemental Material
We present the implementation of the three-body force, proof of the positivity, and the details of the Coulomb interaction. Note that these contents can also be found in other literature; we include it for self-sufficiency. We also discuss details of the optimization algorithm, computational complexity, and finite volume effects. We provide results for nucleon-nucleon phase shifts, odd-even staggering in oxygen isotopes, the pure neutron matter equation of state, and charge density distributions. The numerical results and original data are provided as detailed tables and combined figures.
Auxiliary field transformation for three-body force
For non-perturbative simulations of many-body forces, we employ a discrete auxiliary field that simultaneously simulates two-, three-, and four-body forces without sign oscillations. This transformation has been used in Ref. PLB797-134863* :
$:\exp\left(-\frac{1}{2}C_{2}a_{t}\rho^{2}-\frac{1}{6}C_{3}a_{t}\rho^{3}-\frac{1}{24}C_{4}a_{t}\rho^{4}\right):=\sum_{k=1}^{N}\omega_{k}:\exp\left(\sqrt{-C_{2}a_{t}}\phi_{k}\rho\right):,$ (10)
where $\rho$ is a general one-body density operator, and $C_{2}$, $C_{3}$, $C_{4}$ are the two-, three-, and four-body coefficients, respectively. We solve for the real numbers $\omega_{k}$ and $\phi_{k}$. In this work, we consider only attractive two-body interactions with $C_{2}<0$. To avoid the sign problem, we further require $\omega_{k}>0$ for all $k$.
To determine $\phi_{k}$ and $\omega_{k}$, we expand Eq. (10) up to $\mathcal{O}\left(\rho^{4}\right)$ and compare both sides order by order. For nuclear forces where three- and four-body interactions are typically much weaker than the two-body interaction, we use the ansatz with $N=3$:
$\omega_{1}=\frac{1}{\phi_{1}\left(\phi_{1}-\phi_{3}\right)},\quad\omega_{2}=1+\frac{1}{\phi_{1}\phi_{3}},\quad\omega_{3}=\frac{1}{\phi_{3}\left(\phi_{3}-\phi_{1}\right)},$ (11)
where $\phi_{2}=0$, and $\phi_{1}$ and $\phi_{3}$ are roots of the quadratic equation:
$\phi^{2}+\frac{C_{3}}{\sqrt{-C_{2}^{3}a_{t}}}\phi-\frac{C_{3}^{2}}{C_{2}^{3}a_{t}}+\frac{C_{4}}{C_{2}^{2}a_{t}}-3=0.$ (12)
Using Vieta’s formulas relating polynomial coefficients to sums and products of roots, we verify that Eq. (11) satisfies Eq. (10) up to $\mathcal{O}\left(\rho^{4}\right)$. For a pure two-body interaction ($C_{3,4}=0$), the solution simplifies to $\phi_{1}=-\phi_{3}=\sqrt{3}$, $\phi_{2}=0$, $\omega_{1}=\omega_{3}=1/6$, $\omega_{2}=2/3$. The formalism in Eq. (11) efficiently simulates many-body forces.
In practical calculations, we take $C_{4}=0$ and substitute
$\rho=\overline{\rho}+\frac{C_{s}}{2C_{2}}\overline{\rho}_{s}$ (13)
into Eq. (10). The coefficients $\omega_{k}$ and $\phi_{k}$ are determined by $C_{2}$, $C_{3}$ and $a_{t}$ at the outset. Rather than updating continuous auxiliary fields, we randomly select indices $1\leq k\leq 3$ per lattice site and apply the Metropolis algorithm to accept or reject the new index field. After sufficient warm-up, thermal equilibrium is achieved, and observables are measured by inserting operators at the central time step. Energy is obtained by directly measuring the Hamiltonian expectation value. Note that Eq. (10) generates additional interactions of order $\mathcal{O}(C_{s}^{2}/C_{2})$ or $\mathcal{O}(C_{3}C_{s}/C_{2})$, which are highly suppressed and negligible.
Proof for the positivity of the amplitude
The sufficient condition for a positive definite correlation matrix was given in Ref. PRB71-155115* . Here we present a step-by-step derivation specific to NLEFT calculations. The expectation value of any operator $O$ is expressed as
$\langle O\rangle$ $=$ $\lim_{\tau\rightarrow\infty}\frac{\langle\Phi_{T}|e^{-\tau H/2}Oe^{-\tau H/2}|\Phi_{T}\rangle}{\langle\Phi_{T}|e^{-\tau H}|\Phi_{T}\rangle}=\lim_{L_{t}\rightarrow\infty}\frac{\langle\Phi_{T}|M^{L_{t}/2}OM^{L_{t}/2}|\Phi_{T}\rangle}{\langle\Phi_{T}|M^{L_{t}}|\Phi_{T}\rangle}$
$=$ $\lim_{L_{t}\rightarrow\infty}\frac{\int\mathcal{D}c\exp(-\sum_{i=1}^{L_{t}}c_{i}^{2}/2)\langle\Phi_{T}|M(c_{L_{t}})\cdots M(c_{L_{t}/2+1})OM(c_{L_{t}/2})\cdots M(c_{1})|\Phi_{T}\rangle}{\int\mathcal{D}c\exp(-\sum_{i=1}^{L_{t}}c_{i}^{2}/2)\langle\Phi_{T}|M(c_{L_{t}})\cdots M(c_{1})|\Phi_{T}\rangle},$
where we have neglected the three-body force and applied the traditional Hubbard-Stratonovich auxiliary field transformation and $\tau=L_{t}\ a_{t}$. The proof remains essentially unchanged if we include the three-body force using the discrete transformation. Here $M={:}\exp(-a_{t}H){:}$ is the transfer matrix, $c_{i}(\bm{n})$ is the auxiliary field at the $i$-th time slice, and
$M(c)=\exp\left[\sum_{\bm{n},\bm{n}^{\prime}}a_{t}\frac{\Psi^{\dagger}(\bm{n})\nabla^{2}_{\bm{n}\bm{n}^{\prime}}\Psi(\bm{n}^{\prime})}{2M}+\sum_{\bm{n}}\sqrt{-a_{t}C_{2}}c(\bm{n})\left(\overline{\rho}(\bm{n})+\frac{C_{s}}{2C_{2}}\overline{\rho}_{s}(\bm{n})\right)\right]$ (15)
is the transformed transfer matrix. The summations over $\bm{n}$, $\bm{n}^{\prime}$ run over all $L^{3}$ lattice sites, and $\nabla^{2}_{\bm{n}\bm{n}^{\prime}}$ represents the discretized Laplacian operator implemented using FFT.
The transfer matrix in Eq. (15) is bilinear in the fermion creation and annihilation operators. When acting on a single-particle wave function, it corresponds to a one-body imaginary-time evolution operator of dimension $L^{3}\times L^{3}$. For products of single-particle wave functions, it induces no many-body correlations and evolves each particle independently:
$M(c)\left(|\psi_{1}\rangle\wedge|\psi_{2}\rangle\wedge\cdots\wedge|\psi_{A}\rangle\right)=M(c)|\psi_{1}\rangle\wedge M(c)|\psi_{2}\rangle\wedge\cdots\wedge M(c)|\psi_{A}\rangle,$ (16)
where $\psi_{1,\ldots,A}$ are a set of single-particle wave functions, and $\wedge$ denotes the antisymmetrized product, i.e., the Slater determinant.
Without loss of generality, we assume that $|\Phi_{T}\rangle$ is a Slater determinant composed of single-particle wave functions $\phi_{1},\phi_{2},\ldots,\phi_{A}$. Repeated application of Eq. (16) allows us to rewrite the denominator of Eq. (Proof for the positivity of the amplitude) as
$\mathcal{Z}=\int\mathcal{D}c\exp(-\sum_{i=1}^{L_{t}}c_{i}^{2}/2)\left[\bigwedge_{i=1}^{A}\langle\phi_{i}|\right]\left[\bigwedge_{i=1}^{A}\overline{M}(c)|\phi_{i}\rangle\right]=\int\mathcal{D}c\exp(-\sum_{i=1}^{L_{t}}c_{i}^{2}/2)\det(Z(c)),$ (17)
where $\overline{M}(c)=M(c_{L_{t}})\cdots M(c_{1})$ is the product of single-particle evolution operators, and $Z$ is an $A\times A$ fermionic correlation matrix:
$Z(c)$ $=\bigl{[}\,\langle\phi_{i}\mid\overline{M}(c)\mid\phi_{j}\rangle\bigr{]}_{i,j=1}^{A}=\begin{pmatrix}\langle\phi_{1}\mid\overline{M}(c)\mid\phi_{1}\rangle&\langle\phi_{1}\mid\overline{M}(c)\mid\phi_{2}\rangle&\cdots&\langle\phi_{1}\mid\overline{M}(c)\mid\phi_{A}\rangle\\ \langle\phi_{2}\mid\overline{M}(c)\mid\phi_{1}\rangle&\langle\phi_{2}\mid\overline{M}(c)\mid\phi_{2}\rangle&\cdots&\langle\phi_{2}\mid\overline{M}(c)\mid\phi_{A}\rangle\\ \vdots&\vdots&\ddots&\vdots\\ \langle\phi_{A}\mid\overline{M}(c)\mid\phi_{1}\rangle&\langle\phi_{A}\mid\overline{M}(c)\mid\phi_{2}\rangle&\cdots&\langle\phi_{A}\mid\overline{M}(c)\mid\phi_{A}\rangle\end{pmatrix}.$ (18)
In NLEFT calculations, we generate an ensemble of auxiliary-field configurations $\{c\}$ with the probability distribution
$P(c)\propto\exp(-\sum_{i=1}^{L_{t}}c_{i}^{2}/2)\left|\det(Z(c))\right|$ (19)
and measure observables as arithmetic averages. The sign problem occurs when the determinant in Eq. (19) is not positive definite. In this case, the averaged sign
$\langle e^{i\theta}\rangle=\int\mathcal{D}cP(c)\frac{\det(Z(c))}{\left|\det(Z(c))\right|}$ (20)
quantifies the severity of the sign problem. For $\langle e^{i\theta}\rangle=1$, there is no sign problem, while $\langle e^{i\theta}\rangle\approx 0$ indicates a severe sign problem.
The single nucleon wave functions include spin and isospin indices and can be written as complex vectors of dimension $4\times L^{3}$. Correspondingly, the transfer matrix in Eq. (15) becomes a $4L^{3}\times 4L^{3}$ matrix for any given $c$-field configuration. Since the interaction used in this work has no isospin-mixing term, the determinant factorizes:
$\det(Z(c))=\det(Z_{P}(c))\det(Z_{N}(c)),$ (21)
where $Z_{P}$ and $Z_{N}$ are the correlation matrices for protons and neutrons, respectively. In what follows, we omit the isospin degrees of freedom; all conclusions hold for $Z(c)$, $Z_{P}(c)$, and $Z_{N}(c)$.
For even-even nuclei, we prepare the single nucleon wave functions as paired by the time-reversal operation:
$|\phi_{A/2+k}\rangle=\mathcal{T}|\phi_{k}\rangle,\qquad k=1,2,\ldots,A/2,$ (22)
where $\mathcal{T}=i\sigma_{y}\mathcal{K}$ is the time-reversal operator and $\mathcal{K}$ denotes complex conjugation. The matrix $\overline{M}(c)$ commutes with $\mathcal{T}$, as the spin Pauli matrices and the imaginary factor $i$ in $\rho_{s}$ both contribute sign changes that cancel under time reversal. Note that we focus on the $\mathcal{T}$-symmetry of the decomposed transfer matrix in Eq. (15), not the original interactions, which correspond to squares of the decomposed interactions and are always time-reversal even. For example, while the Coulomb force is time-reversal invariant, its auxiliary field transformation explicitly involves an imaginary phase and cannot be simulated without a sign problem.
Using the properties of $\mathcal{T}$ and Eq. (22), we obtain the relations:
$\langle\phi_{A/2+i}|\overline{M}(c)|\phi_{A/2+j}\rangle$ $=$ $\langle\mathcal{T}\phi_{i}|\overline{M}(c)|\mathcal{T}\phi_{j}\rangle=\langle\phi_{i}|\overline{M}(c)|\mathcal{T}^{\dagger}\mathcal{T}\phi_{j}\rangle^{*}=\langle\phi_{i}|\overline{M}(c)|\phi_{j}\rangle^{*},$ (23)
$\langle\phi_{A/2+i}|\overline{M}(c)|\phi_{j}\rangle$ $=$ $\langle\mathcal{T}\phi_{i}|\overline{M}(c)|\phi_{j}\rangle=\langle\phi_{i}|\overline{M}(c)|\mathcal{T}^{\dagger}\phi_{j}\rangle^{*}=-\langle\phi_{i}|\overline{M}(c)|\phi_{A/2+j}\rangle^{*}.$
The correlation matrix then has the structure:
$Z\;=\;\begin{pmatrix}U&-V^{*}\\[4.0pt] V&U^{*}\end{pmatrix},$ (24)
where $U$ and $V$ are $A/2\times A/2$ complex matrices.
We define a spin-flipping matrix:
$\Sigma\;=\;i\sigma_{y}\otimes I_{A/2}\;=\;\begin{pmatrix}0&I_{A/2}\\[6.0pt] -I_{A/2}&0\end{pmatrix},$ (25)
where $I_{A/2}$ is the $A/2\times A/2$ identity matrix. Direct verification shows that
$Z\,\Sigma\;=\;\Sigma\,Z^{*}.$ (26)
For any eigenvalue $\lambda\in\mathbb{C}$ with corresponding eigenvector $v$ satisfying $Z\,v=\lambda\,v$, we have
$Z\,(\Sigma\,v^{*})=\Sigma\,Z^{*}\,v^{*}=\Sigma\,(\lambda^{*}\,v^{*})=\lambda^{*}\,(\Sigma\,v^{*}),$ (27)
implying that $\lambda^{*}$ is also an eigenvalue with eigenvector $w=\Sigma v^{*}$. Thus, complex eigenvalues of $Z$ always appear in conjugate pairs. If $\lambda$ is real, we find
$\langle v\mid w\rangle=v^{\dagger}\,\Sigma\,v^{*}=(v^{T}\,\Sigma\,v)^{*}=0,$ (28)
indicating that $v$ and $w$ are orthogonal eigenvectors corresponding to the same real eigenvalue. In this case, the real eigenvalue appears twice in the diagonalized form of $Z$. Consequently, the determinant $\det(Z)$, being the product of all eigenvalues, is always nonnegative.
Coulomb force
A static Coulomb force $V_{\mathrm{cou}}$ is included. The form is the same as that used in Ref. EPJA61-85* with a different cutoff. We compute the first-order perturbative energy of $V_{\mathrm{cou}}$. In momentum space, we have
$V_{\mathrm{cou}}=\frac{\alpha}{q^{2}}\exp\!\Bigl{(}-\frac{q^{2}}{2\Lambda_{\mathrm{cou}}^{2}}\Bigr{)},$ (29)
where $\alpha=1/137$ is the fine structure constant, $\Lambda_{\mathrm{cou}}=180\,\mathrm{MeV}$ is a momentum cutoff that removes the singularity at $r=0$, and $\bm{q}=\bm{p}^{\prime}-\bm{p}$ is the momentum transfer. In coordinate space, the potential is expressed using the error function:
$V_{\mathrm{cou}}(r)=\alpha\,\frac{\mathrm{erf}(\Lambda_{\mathrm{cou}}\,r/2)}{r}.$ (30)
The asymptotic behavior satisfies
$V_{\mathrm{cou}}(0)=\frac{\alpha\,\Lambda_{\mathrm{cou}}}{\sqrt{\pi}}\quad\text{and}\quad V_{\mathrm{cou}}(r)\to\frac{\alpha}{r}\quad\text{as}\;\Lambda_{\mathrm{cou}}\to\infty.$
The potential $V_{\mathrm{cou}}$ is smooth for all $r$ and can be implemented on the lattice using FFT.
Gradient Descent Algorithm for parameter optimization
Gradient descent is an optimization algorithm that leverages gradient information to iteratively update parameters. Given the computational intensity and statistical uncertainties inherent in QMC calculations, complex optimization schemes are infeasible. We adopt a stable and efficient gradient descent method to perform the optimization task, summarized as follows:
  1. 1.
    Initialization: Select a reasonable initial parameter set $\theta_{0}$.
  2. 2.
    Parameter rescaling: Rescale parameters to ensure comparable variation scales. During early iterations:
    $\theta_{i}\rightarrow\theta_{i}\left|\frac{\partial\chi^{2}}{\partial\theta_{i}}\right|,$
    yielding partial derivatives of magnitude $\sim 1$. Near convergence when iterations slow, apply fixed scaling:
    $\theta_{i}\rightarrow\theta_{i}\times k_{i},$
    with scaling factors: $k_{s_{L}}=1.853\times 10^{5},k_{C_{2}}=3.946\times 10^{11},k_{C_{3}}=1.004\times 10^{18},$ and $k_{C_{s}}=1.260\times 10^{15}.$ This technique significantly accelerates convergence.
  3. 3.
    Gradient computation: At iteration $t$, compute partial derivatives of the loss function. For $C_{2}$, $C_{3}$, and $C_{s}$, directly measure energy contributions from corresponding terms. For $s_{\rm L}$ (nonlinear in binding energy), apply the Feynman-Hellmann theorem for derivative estimation, achieving Monte Carlo uncertainties $<0.1\%$.
  4. 4.
    Parameter update:
    $\theta_{t+1}=\theta_{t}-\alpha\frac{\partial\chi^{2}(\theta_{t})}{\partial\theta},$
    with adaptive learning rate:
    $\alpha=\begin{cases}\alpha\times 2&\text{if }\chi^{2}\text{ decreases}\\ \alpha\times 0.2&\text{otherwise}\end{cases}.$
    This strategy increases step size in smooth regions and decreases it to prevent instability or overshooting.
  5. 5.
    Convergence check: Terminate when:
    $|\chi^{2}(\theta_{t+1})-\chi^{2}(\theta_{t})|<\epsilon.$
Figure 5: optimization of loss function. (Upper) Loss function $\chi^{2}$ versus iteration steps. (Lower) Dimensionless parameters normalized to converged values. Deviations of scaled $s_{\rm L}$ and $C_{2}$ are exaggerated for visibility.
Fig. 5 (Upper) demonstrates the monotonic decrease of $\chi^{2}$ during optimization, confirming the method’s efficacy. Fig. 5 (Lower) schematically illustrates parameter convergence to final values.
Computational complexity for heavy nuclei
The sign-problem-free QMC method is particularly suited for simulating heavy nuclei, as its computational time scales polynomially with nucleon number, whereas the sign problem inevitably introduces exponential scaling. We numerically demonstrate this advantage below.
Fig. 6(a) depicts the average CPU hours per auxiliary-field configuration, measured on a consistent supercomputing platform. The dashed parabolic curve serves as a visual guide. Using fixed parameters ($L=11$, $L_{t}=200$), we confirm polynomial time complexity and observe approximately linear scaling for $A\geq 40$.
Fig. 6(b) displays statistical uncertainties in binding energies across 16,800 configurations. Uncertainties increase gradually with mass number but remain below 1 MeV for nuclei up to $A\sim 100$. Relative statistical errors (Fig. 6(d)) exhibit rapid convergence in heavy nuclei, stabilizing near one-thousandth for $A\geq 40$.
Combining these observations, we achieve this accuracy level with linearly increasing computational resources. This scalability is highly encouraging for QMC studies of heavy nuclei and warrants further investigation. Although larger simulation boxes are required for $A>100$ to suppress finite-volume effects, high-precision calculations remain feasible.
Finally, Fig. 6(c) compares the average sign $\langle e^{i\theta}\rangle$ between our sign-problem-free interaction and a non-perturbative N2LO chiral force. The chiral interaction, featuring complex spin/isospin dependencies, exhibits a sign problem where $\langle e^{i\theta}\rangle$ decreases rapidly with mass number. Conversely, our approach maintains $\langle e^{i\theta}\rangle=1$ for all systems.
Figure 6: Computational performance. (a) Average CPU hours per configuration; (b) Absolute binding energy statistical uncertainties (16,800 configurations); (c) Average sign $\langle e^{i\theta}\rangle$ for LAT-OPT1 (orange squares) versus N2LO chiral force (blue circles, sign-problematic); (d) Relative binding energy statistical uncertainties.
Finite Volume Effects
All calculations in this work employ a lattice box of size $L=11$, which yields negligible finite-volume effects for binding-energy computations of most nuclei. This is illustrated in Fig. 7, showing convergence of binding energies (excluding Coulomb interactions) for ${}^{40}\mathrm{Ca}$ ($L_{t}=400$) and ${}^{100}\mathrm{Sn}$ ($L_{t}=200$) as box size increases. The distinct plateau beyond $L=10$ confirms effective convergence of finite-volume effects.
Figure 7: Finite-volume convergence of binding energies. Binding energies (Coulomb excluded) for ${}^{40}\mathrm{Ca}$ ($L_{t}=400$, left) and ${}^{100}\mathrm{Sn}$ ($L_{t}=200$, right) versus lattice size $L$. Blue symbols: Monte Carlo results with statistical uncertainties (error bars); red curves: exponential fits $E(\infty)+Ce^{-kL}$. Plateau formation at $L\gtrsim 10$ confirms negligible finite-volume effects.
Nucleon-nucleon scattering phase shifts
Nuclear ab initio calculations aim to solve nuclear structure from bare nuclear forces constrained solely by nucleon-nucleon phase shifts and few-body data. However, reconciling theoretical predictions with experimental data remains challenging. Consequently, modern ab initio approaches often employ interactions constrained by both NN phase shifts and nuclear properties.
We therefore investigate the inverse problem: whether nuclear forces determined solely by finite-nuclei observables can reproduce NN phase shifts. This approach is theoretically feasible since nuclear masses and charge radii can be measured with higher precision than scattering cross sections. Conversely, these observables should encode complete information about nuclear forces. Solving such inverse many-body problems requires high-precision nuclear solvers. While numerous mean-field and density functional interactions are calibrated to finite nuclei, these methods are based on variational principles and not exact. The sign-problem-free QMC method provides the essential foundation for this problem, offering both an unbiased high-precision solver and an efficient optimizer for nuclear masses.
Fig. 8 shows that our LAT-OPT1 predictions for NN S-wave phase shifts fall between empirical ${}^{1}S_{0}$ and ${}^{3}S_{1}$ values across the momentum range up to $250\,\mathrm{MeV}$, demonstrating reasonable reproduction of S-wave interaction strength. Fitting the phase shifts to the effective-range expansion,
$k\cot\delta=-\frac{1}{a_{0}}+\frac{1}{2}\,r_{0}\,k^{2}+\cdots,$
yields a scattering length $a_{0}=6.86\,\mathrm{fm}$ and effective range $r_{0}=2.10\,\mathrm{fm}$, both within the range of ${}^{1}S_{0}$ and ${}^{3}S_{1}$ experimental values. This represents a leading-order phase shift prediction from nuclear structure data.
While this simple interaction cannot reproduce S-wave splitting, S-D mixing, or P-wave phase shifts, including additional spin/isospin-dependent operators and pion-exchange potentials should improve description accuracy. Such extensions would however require sign-problem mitigation techniques.
Figure 8: S-wave nucleon-nucleon phase shifts. Empirical phase shifts $\delta$ for ${}^{1}S_{0}$ (blue circles) and ${}^{3}S_{1}$ (red squares) channels versus relative momentum $P_{\rm rel}$ PhysRevC.48.792*. Black dash-dot curve: LAT-OPT1 predictions. Inset: Extracted scattering length $a_{0}$ and effective range $r_{0}$. LAT-OPT1 accurately captures the empirical trend up to $P_{\rm rel}\approx 250\,\mathrm{MeV}$.
Oxygen isotopes and pure neutron matter
Fig. 9 presents binding energies per nucleon across the oxygen isotopic chain for both even-$A$ and odd-$A$ nuclei. The inset displays corresponding average phases. Calculations for odd nucleon numbers introduce a mild sign oscillation due to slight time-reversal symmetry breaking, yet yield statistical uncertainties comparable to even-$A$ nuclei, indicating no significant efficiency degradation. The computed energies reproduce odd-even staggering with slightly overestimated pairing gaps. Discrepancies at large neutron numbers may indicate missing interaction components (e.g., tensor forces).
Figure 9: Binding energies in oxygen isotopes. Binding energy per nucleon $E/A$ versus mass number $A$. Blue circles: even-$A$; red circles: odd-$A$; black crosses: experimental values. Inset: Average phase $\langle e^{i\theta}\rangle$ ($=1$: sign-problem-free).
Fig. 10 compares pure neutron matter equations of state predicted by different interaction models. The LAT-OPT1 parameterization shows excellent agreement with benchmark results up to saturation density. As demonstrated in the main text, this model approximately reproduces empirical saturation properties of symmetric nuclear matter. These results confirm that infinite nuclear matter properties can be reliably extracted from light nuclei fitting ($A\leq 40$), suggesting the essential role of many-body correlations.
Figure 10: Pure neutron matter equation of state. Energy per particle versus neutron density for LAT-OPT1 interaction (red squares). Comparisons include: N3LO chiral interactions (EM 500 MeV, EGM 450/500 MeV, EGM 450/700 MeV) Tews2013* ; variational APR Akmal1998* ; Auxiliary Field Diffusion MC (GCR) Gandolfi2012* .
Charge density distributions and charge radii
Charge density distributions (Fig. 11) and corresponding charge radii (Table 2) for the six target nuclei were calculated using the pinhole algorithm Elhatisari2017* . This method induces sign problems and is typically restricted to light nuclei ($A\leq 40$). Calculations incorporated the proton size $R_{p}=0.84\,\mathrm{fm}$ via convolution.
Results with and without perturbative Coulomb inclusion demonstrate only minor corrections for these light nuclei. The computed charge densities approximately reproduce empirical profiles, though central regions exhibit large statistical uncertainties and slight underestimation compared to experimental data. Charge radii are overestimated by $\sim 5$-$7\%$. Inclusion of charge radii as optimization constraints should improve these results.
Figure 11: Charge density distributions. Calculated charge density distributions for selected nuclei without (blue circles) and with (red squares) perturbative Coulomb corrections compared with the empirical results DeVries1987* .
Comparison with phenomenological models
When evaluated as a nuclear mass model, LAT-OPT1 predictions compare favorably with established approaches: macroscopic-microscopic (mac-mic) models, relativistic mean field (RMF) theory, and Skyrme density functional theory. These models demonstrate strong performance in describing nuclear structures, with parameters constrained by global fitting to binding energies, charge radii, and nuclear matter saturation properties across training sets typically spanning 16O to 208Pb.
We compare three published parameterizations: the FRDM Moller2016* (mac-mic), PC-PK1 Zhang2022* (RMF), and UNEDF1 PRC85-024304* (Skyrme). Standard deviations from experimental masses are computed for even-even nuclei with $N,Z\leq 28$ (Table 3). Phenomenological models show $\sim 2$-$3$ MeV RMS deviations in this region, improving significantly for heavier nuclei where mean-field contributions dominate. LAT-OPT1 achieves comparable precision ($\sim 3$ MeV RMS) despite fewer parameters, while mac-mic models yield smaller deviations with more adjustable parameters.
Table 3 also includes results from the Wigner-SU(4) interaction in Ref. PLB797-134863* , demonstrating significant improvement through: (1) systematic finite-volume effect elimination via larger boxes; (2) comprehensive imaginary-time extrapolation replacing fixed-$\tau$ calculations. The inclusion of spin-orbit coupling and efficient optimization further reduced deviations from $\sim 10$ MeV to $3$ MeV. Future extensions to larger training sets may further improve accuracy.
Complete numerical results (Table LABEL:tab:gse) compare LAT-OPT1 with mean-field calculations, listing Coulomb contributions separately. We also present the original data as separate figures for each nuclei. We show the results without the Coulomb energy, which can be restored by looking up Table LABEL:tab:gse.
Table 2: Charge radii. RMS charge radii $R$ (fm) for selected nuclei computed with LAT-OPT1 versus experimental values exp* .
Nucleus 4He 16O 24Mg 28Si 32S 40Ca
LAT-OPT1 1.797(3) 2.872(6) 3.235(24) 3.319(25) 3.409(30) 3.537(23)
Experiment 1.6755(28) 2.6991(52) 3.0570(16) 3.1224(24) 3.2611(18) 3.4776(19)
Relative error [%] 7.3 6.4 5.8 6.3 4.5 1.7
Table 3: Standard deviations $\sigma$ of ground-state binding energies for even-even nuclei ($N,Z\leq 28$) across mass models. Key: "Parameters" = adjustable parameter count; "Nuclei" = nuclei count for $\sigma$; $\sigma_{\rm all}$ = standard deviation over full nuclide chart.
Model Parameterization Parameters $\sigma$ (MeV) Nuclei $\sigma_{\rm all}$ (MeV)
Macroscopic-microscopic FRDM Moller2016* $>30$ 1.15 69 0.56
Relativistic mean field PC-PK1 Zhang2022* 11 2.25 60 1.52
Skyrme DFT UNEDF1 PRC85-024304* 12 3.43 75 1.91
Lattice EFT Wigner-SU(4) PLB797-134863* 4 10.21 55
Lattice EFT LAT-OPT1 5 2.93 76
Table 4: Ground-state binding energies of even-even nuclei calculated with UNEDF1 PRC85-024304* , PC-PK1 Zhang2022* and LAT-OPT1, alongside experimental values (EXP) and perturbative Coulomb corrections in LAT-OPT1 calculations (COU). The LAT-OPT1 errors are combined statistical and imaginary-time-extrapolation errors. All energies are in MeV.
$A$ $N$ UNEDF1 PC-PK1 LAT-OPT1 EXP COU (MeV)
Z = 2 (He)
4 2 --18.70 --28.63 --28.30 0.543
6 4 --23.24 --33.17(6) --29.27 0.527
8 6 --24.61 --37.68(10) --31.39 0.543
Z = 4 (Be)
8 4 --43.42 --56.98(20) --56.50 2.450
10 6 --58.52 --66.96(16) --64.97 2.533
12 8 --64.89 --72.29(30) --68.64 2.560
14 10 --77.28(29) --69.916 2.482
Z = 6 (C)
10 4 --54.07 --63.45(16) --60.32 6.056
12 6 --83.11 --91.07(12) --92.16 6.180
14 8 --99.80 --104.60(10) --105.28 6.297
16 10 --105.23 --112.88(30) --110.75 6.175
18 12 --109.27 --119.75(51) --115.67 6.038
20 14 --112.72 --125.61(47) --119.22 6.047
22 16 --114.57 --131.86(55) --119.26 6.001
Z = 8 (O)
12 4 --55.10 --59.70 --63.70(12) --58.58 10.875
14 6 --93.98 --101.16 --99.62(10) --98.73 11.278
16 8 --121.71 --128.03 --126.94(15) --127.62 11.318
18 10 --136.12 --141.84 --139.33(22) --139.81 11.221
20 12 --147.62 --152.98 --152.22(21) --151.37 11.142
22 14 --157.22 --161.39 --163.63(26) --162.03 11.099
24 16 --164.00 --165.98 --174.87(30) --168.95 11.051
26 18 --165.45 --167.41 --179.02(34) --168.93 11.873
Z = 10 (Ne)
16 6 --100.15 --101.86(16) --97.33 17.054
18 8 --129.33 --134.37 --133.48(40) --132.14 17.190
20 10 --152.93 --160.81 --158.90(52) --160.64 17.123
22 12 --172.32 --177.97 --176.33(25) --177.77 17.011
24 14 --186.86 --192.61 --190.22(23) --191.84 17.112
26 16 --198.32 --202.25 --203.86(16) --201.55 17.052
28 18 --204.30 --208.55 --211.98(36) --206.87 16.867
30 20 --207.95 --212.44 --217.41(29) --211.04 16.653
Z = 12 (Mg)
20 8 --132.57 --135.54 --138.08(20) --134.56 24.448
22 10 --163.81 --169.09 --168.72(22) --168.58 24.334
24 12 --190.53 --206.16 --196.03(32) --198.26 24.424
26 14 --211.19 --220.52 --213.13(32) --216.68 24.379
28 16 --227.37 --231.59 --229.96(27) --231.63 24.412
30 18 --238.48 --243.12 --242.85(18) --241.63 24.193
32 20 --246.65 --252.21 --250.49(21) --249.72 23.985
34 22 --251.27 --260.61 --259.95(36) --256.71 23.616
36 24 --255.81 --266.85 --267.76(33) --260.80 23.399
Z = 14 (Si)
24 10 --168.73 --172.61 --174.45(41) --172.01 32.929
26 12 --201.15 --207.64 --204.98(50) --206.04 32.943
28 14 --229.40 --239.37 --233.18(35) --236.54 33.227
30 16 --251.31 --253.02 --252.60(21) --255.62 33.015
32 18 --267.55 --271.61 --268.49(36) --271.41 32.872
34 20 --280.56 --284.74 --281.45(27) --283.46 32.687
36 22 --288.46 --292.93 --292.97(39) --292.05 32.318
38 24 --295.95 --303.93 --302.81(39) --299.93 31.932
40 26 --301.71 --311.49 --311.39(76) --306.23 31.677
Z = 16 (S)
28 12 --206.45 --209.32 --212.12(28) --209.41 42.482
30 14 --240.24 --241.38 --242.86(21) --243.68 42.741
32 16 --267.57 --269.35 --271.34(32) --271.78 42.930
34 18 --288.88 --289.29 --291.99(11) --291.84 42.446
36 20 --306.68 --308.37 --308.54(20) --308.71 42.294
38 22 --318.83 --322.56 --321.67(26) --321.05 41.941
40 24 --330.10 --335.55 --334.04(77) --333.17 41.665
42 26 --339.61 --346.39 --345.48(43) --344.12 41.420
44 28 --347.49 --354.35 --354.82(54) --351.82 41.284
Z = 18 (Ar)
32 14 --244.70 --246.95 --248.05(29) --246.40 53.414
34 16 --276.88 --276.40 --281.05(16) --278.72 53.397
36 18 --303.51 --306.28 --306.11(40) --306.72 53.459
38 20 --326.49 --327.13 --328.09(14) --327.34 53.073
40 22 --343.13 --343.28 --344.37(28) --343.81 52.717
42 24 --357.79 --360.56 --360.18(32) --359.34 52.351
44 26 --371.15 --374.98 --373.95(46) --373.73 52.170
46 28 --382.74 --387.54 --387.62(32) --386.97 52.050
Z = 20 (Ca)
36 16 --281.41 --281.45 --285.52(20) --281.37 65.306
38 18 --313.11 --313.03 --315.74(10) --313.12 65.166
40 20 --341.11 --343.06 --343.00(18) --342.05 65.436
42 22 --362.24 --363.51 --362.55(21) --361.90 64.697
44 24 --381.13 --382.12 --380.36(33) --380.96 64.434
46 26 --398.49 --399.38 --397.78(12) --398.77 64.181
48 28 --413.75 --415.50 --414.49(29) --416.00 64.110
Z = 22 (Ti)
40 18 --315.78 --314.46 --319.86(38) --314.63 77.408
42 20 --348.27 --348.76 --349.70(17) --346.89 77.381
44 22 --373.46 --373.09 --376.20(45) --375.47 77.218
46 24 --396.43 --398.46 --396.45(44) --398.20 76.715
48 26 --417.47 --419.46 --416.18(28) --418.70 76.576
50 28 --436.29 --436.72 --434.70(46) --437.79 76.446
Z = 24 (Cr)
44 20 --351.82 --351.51 --354.10(46) --349.78 90.962
46 22 --381.05 --382.51 --383.20(76) --381.98 90.616
48 24 --409.68 --412.29 --411.70(59) --411.47 90.573
50 26 --434.00 --435.67 --431.81(20) --435.05 90.115
52 28 --455.12 --454.84 --451.75(32) --456.35 90.030
Z = 26 (Fe)
48 22 --385.80 --386.47 --386.93(27) --385.09 105.36
50 24 --417.55 --418.55 --417.94(71) --417.70 105.05
52 26 --445.60 --447.92 --447.08(78) --447.70 105.25
54 28 --470.35 --470.35 --467.81(40) --471.76 104.86
Z = 28 (Ni)
52 24 --421.10 --419.22 --421.19(45) --420.36 120.98
54 26 --452.63 --451.87 --450.88(32) --453.22 120.80
56 28 --481.59 --483.65 --479.33(60) --484.00 121.35
Z = 40 (Zr)
80 40 --669.13 --666.35 --672.1(8) --669.20 233.90
90 50 --786.02 --783.47 --782.1(5) --783.90 230.52
Z = 50 (Sn)
100 50 --828.83 --827.97 --824.7(8) --825.16 348.76
120 70 --1020.24 --1019.77 --1033.65(11) --1020.48 339.35
132 82 --1101.53 --1104.47 --1134.7(27) --1102.84 333.14